# LANTERN++: Enhanced Relaxed Speculative Decoding with Static Tree Drafting for Visual Auto-regressive Models

### v1

@@ Abstract
Speculative decoding has been widely used to accelerate autoregressive (AR) text generation. However, its effectiveness in visual AR models remains limited due to token selection ambiguity, where multiple tokens receive similarly low probabilities, reducing acceptance rates. While dynamic tree drafting has been proposed to improve speculative decoding, we show that it fails to mitigate token selection ambiguity, resulting in shallow draft trees and suboptimal acceleration. To address this, we introduce LANTERN++, a novel framework that integrates static tree drafting with a relaxed acceptance condition, allowing drafts to be selected independently of low-confidence predictions. This enables deeper accepted sequences, improving decoding efficiency while preserving image quality. Extensive experiments on state-of-the-art visual AR models demonstrate that LANTERN++ significantly accelerates inference, achieving up to ×$\mathbf{\times 2.56}$ speedup over standard AR decoding while maintaining high image quality.
@@ 1 Introduction
Recent advances in speculative decoding have significantly accelerated auto-regressive (AR) generation in language models (Leviathan et al., 2023; Cai et al., 2024; Li et al., 2024a; b). Extending this approach to visual AR models (Sun et al., 2024; Team, 2024; Chern et al., 2024; Liu et al., 2024) presents unique challenges due to the inherent nature of image token distributions. In particular, Jang et al. (2025) identified a key issue, token selection ambiguity, wherein multiple next-tokens receive nearly equivalent probabilities. This results in the drafter model struggling to accurately predict the next token, since even the target model itself lacks a highly confident choice. Consequently, speculative acceleration becomes ineffective as the drafter’s predictions tend to be unreliable.
LANTERN (Jang et al., 2025) addressed this by introducing a relaxed acceptance condition leveraging latent-space similarities, where tokens that are close in the latent space are treated as visually similar. This method improved inference speed while maintaining high image quality. However, LANTERN was implemented using EAGLE-2’s (Li et al., 2024b) dynamic tree drafting, which adaptively selects draft tokens based on the drafter’s confidence. In visual AR models, token selection ambiguity leads to two major issues in dynamic tree drafting: (1) low-confidence scores from the drafter result in excessively shallow draft trees, preventing the formation of long accepted sequences, and (2) draft tokens are deterministically selected based on drafter confidence, leading to overly low acceptance probabilities. These two factors collectively degrade the overall efficiency of speculative decoding.
@@
To overcome these limitations, we introduce LANTERN++, an enhanced framework that revises two key aspects: (1) the adoption of static tree drafting and (2) the integration of a multiplicative bound tailored to static tree drafting. By employing static tree drafting, LANTERN++ ensures that deep draft sequences can be generated even under low-confidence conditions. Moreover, static tree drafting allows draft tokens to be selected stochastically rather than deterministically based on confidence, effectively restoring a normal level of acceptance probability. Additionally, LANTERN++ refines the relaxation mechanism by replacing the additive acceptance boost with a multiplicative distributional distortion bound, which scales acceptance probabilities relative to the original distribution. This adjustment prevents disproportionate favoring of certain tokens, ensuring stable probability scaling and preserving distributional consistency. Extensive experiments demonstrate that LANTERN++ achieves substantial speedups while maintaining high image quality, making it a promising direction for efficient visual AR modeling.
@@
Figure 1: Images generated by LANTERN++ on Lumina-mGPT exhibit remarkable acceleration in decoding steps while maintaining high generation quality. The top row shows images generated by standard auto-regressive decoding, while the bottom row displays images generated with acceleration through LANTERN++. The step compression ratio is presented in white at the bottom-right of each image.
@@
  • We analyze dynamic vs. static tree drafting in speculative decoding and show that static tree drafting enables longer accepted sequences by removing confidence-based constraints.
  • We introduce LANTERN++, which integrates static tree drafting with a multiplicative relaxation mechanism to ensure stable and consistent acceptance probabilities.
  • Extensive experiments demonstrate that LANTERN++ achieves up to ×$\times 2.56$ lower latency and ×$\times 3.63$ step compression while maintaining high image quality.
@@ 2 Preliminaries
Tree Drafting and Verification
@@
Unlike standard chain drafting (Leviathan et al., 2023), which generates a single draft token sequence, tree drafting (Cai et al., 2024; Li et al., 2024a; b) samples multiple draft tokens per forward pass, forming a tree structure for diverse draft sequences. Given an input $X_{1:N}$, the drafter first samples $s_{1}$ tokens $\hat{x}_{1,1},\ldots,\hat{x}_{1,s_{1}}$ from $p(x|X_{1:N})$, then expands each token with $s_{2}$ subsequent drafts in a single forward pass using a tree-aware attention mask. This process continues up to depth $K$, constructing a draft tree with only $K$ forward passes. The tree is then flattened and verified by the target model in a single forward pass.
@@
Tree drafting has two main variants. Static tree drafting (Cai et al., 2024; Li et al., 2024a) uses a fixed tree structure where a predefined set of draft tokens are generated at each level. Dynamic tree drafting (Li et al., 2024b) adapts the tree structure based on context. It estimates a global accept probability for each token as $V_{i}=\prod_{t_{j}\in\text{Path}(\text{root},t_{i})}p_{j}\approx\prod_{t_{j}\in\text{Path}(\text{root},t_{i})}c_{j}$, where $p_{j}$ is the actual accept probability and $c_{j}$ is the drafter’s confidence score. The model then expands the top-$k$ tokens and trims the set of drafts by $V_{i}$ before verification.
Token Selection Ambiguity and LANTERN
Visual AR models exhibit token selection ambiguity (Jang et al., 2025), where next-token distributions are more dispersed than in language models (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023), resulting in the drafter often failing to predict the target model’s outputs accurately and thus limiting the overall speedup achievable via speculative decoding. To address this, LANTERN (Jang et al., 2025) exploits latent proximity by treating visually similar (close in latent space) tokens as interchangeable. Given a draft token $\widetilde{x}$, it identifies its $k$ nearest neighbors in latent space, $B_{k}(\widetilde{x})$, and applies a Total Variation Distance (TVD) constraint $\delta>0$ to refine the neighborhood into $A_{k,\delta}(\widetilde{x})$, ensuring minimal distributional distortion. The acceptance condition is then relaxed from $\min(1,\frac{q(\widetilde{x}|s)}{p(\widetilde{x}|s)})$ to $\min\left(1,\sum_{x\in A_{k,\delta}(\widetilde{x})}\frac{q(x|s)}{p(\widetilde{x}|s)}\right)$, boosting the effective acceptance rate while preserving image quality. Additional preliminaries are provided in Appendix A.
@@ 3 Why Dynamic Tree Drafting Fails in Visual AR Models
Applying dynamic tree drafting to visual AR models, where token selection ambiguity is prevalent, introduces two key issues: (1) draft trees become overly shallow due to low drafter confidence, and (2) the deterministic selection of draft tokens results in consistently low acceptance probabilities. These factors fundamentally limit the effectiveness of speculative decoding by restricting the acceptance of longer draft sequences, thereby reducing overall acceleration.
Shallow Draft Tree.
Dynamic tree drafting computes a global accept probability for each draft node in the draft tree as the product of the drafter’s confidence scores along the path from the root: $V_{i}=\prod_{t_{j}\in\text{Path}(\text{root},t_{i})}c_{j}$. However, in visual AR models, where drafter confidence scores tend to be low, this product decays exponentially as draft sequence length increases. Consequently, longer draft sequences become significantly less likely to survive, leaving only shorter ones. The result is a draft tree that is wide (many branches) but shallow, preventing the formation of long accepted sequences. Since longer sequences are key to achieving high step compression, this structural limitation fundamentally restricts the acceleration potential of dynamic tree drafting. An example resulting shallow tree can be found in Appendix C.1.
Low Accept Probabilities.
In speculative sampling (Leviathan et al., 2023), a draft token $\widetilde{x}$ is accepted with probability $\min(1,q(\widetilde{x}|s)/p(\widetilde{x}|s))$, which depends on the alignment between the drafter’s predictions and the target model’s likelihood. However, in dynamic tree drafting, tokens are deterministically selected based on the drafter’s confidence, effectively setting $p(\widetilde{x}|s)=1$. This reduces the acceptance probability to $q(\widetilde{x}|s)$. Due to token selection ambiguity, visual AR models produce highly dispersed next-token distributions, with even the most likely token often receiving a probability below 0.2. As a result, even accurate drafter predictions lead to an unnecessarily low acceptance rate.
@@
Table 1 highlights the contrasting effects of dynamic tree drafting in language models and visual AR models. In language models, dynamic tree drafting improves the step compression ratio from ×$\times 3.94$ to ×$\times 4.98$, leveraging concentrated next-token distributions and high-confidence drafter predictions to construct deeper draft trees. In visual AR models, however, token selection ambiguity leads to dispersed next-token distributions and lower drafter confidence, producing wide yet shallow draft trees. Consequently, static tree drafting achieves a higher step compression ratio (×$\times 2.49$) than dynamic tree drafting (×$\times 2.11$), underscoring the latter’s limitations in handling token selection ambiguity.
@@ 4 LANTERN++: Relaxed Acceptance with Static Tree Drafting
As demonstrated in Section 3, dynamic tree drafting in visual AR models suffers from limited tree depth and low accept probabilities due to token selection ambiguity. Static tree drafting, in contrast, avoids these pitfalls by relying on a fixed tree structure rather than confidence-based path selection. However, it still struggles with low accept probabilities, as visual AR models inherently assign low probabilities to individual tokens. To address this, we introduce LANTERN++, a refined relaxation of the acceptance condition tailored for static tree drafting.
A key challenge in applying relaxation methods to static tree drafting is the variability of $p(x|s)$. Unlike dynamic tree drafting, where draft tokens are deterministically selected ($p(x|s)=1$), static tree drafting maintains the true drafter probabilities, which can vary significantly across tokens. This variability makes the original additive relaxation strategy problematic: when $p(x|s)$ is small, an additive increase $\delta$ can lead to disproportionately high accept probabilities, while for larger $p(x|s)$, the boost becomes negligible. This lack of proportionality results in inconsistent behavior and potential distortion of the target distribution.
To overcome this issue, LANTERN++ replaces the additive relaxation with a multiplicative distributional distortion bound $\lambda$, ensuring that accept probability adjustments remain proportional to the target model’s likelihood. Specifically, for each draft token $\widehat{x}$, its neighborhood $B_{k}(\widehat{x})$ is first determined in the same manner as in LANTERN. A refined subset $A_{k,\lambda}(\widehat{x})\subseteq B_{k}(\widehat{x})$ is then selected such that
$\sum_{x\in A_{k,\lambda}(\widehat{x})}q(x|s)<\lambda\,q(\widehat{x}|s),$
@@
This multiplicative relaxation offers two key advantages. First, it maintains a consistent proportional adjustment across different values of $p(\widehat{x}|s)$, ensuring that the acceptance probability is neither excessively inflated nor suppressed. Second, it prevents instability in cases where $p(\widehat{x}|s)$ is small, as a fixed additive boost could lead to over-amplification. By contrast, the multiplicative bound scales naturally with $q(x|s)$, yielding a more stable and controlled relaxation.
In summary, LANTERN++ refines the acceptance condition for static tree drafting by introducing a multiplicative bound that maintains proportionality and stability. This adjustment leads to a more predictable relaxation mechanism, improved alignment with the target distribution, and better acceleration performance in visual AR models.
@@ 5 Experiments
Table 2: Acceleration performance and image generation performance of standard decoding, EAGLE-1, and LANTERN++ over different visual AR models on MS-COCO 2017 validation captions.
@@
We conduct extensive experiments to evaluate the effectiveness of LANTERN++ in accelerating inference for visual AR models while preserving image quality. Our evaluation compares LANTERN++ against standard autoregressive (AR) decoding and EAGLE-1 (Li et al., 2024a) across multiple state-of-the-art visual AR models, including Lumina-mGPT (Liu et al., 2024), Anole (Chern et al., 2024), and LlamaGen-XL (Sun et al., 2024). All experiments are conducted on the MS-COCO 2017 validation set (Lin et al., 2014) using a single NVIDIA A100 PCIe 80GB GPU. More details on the experimental setting can be found in Appendix B.
Table 2 presents the step compression ratios and inference latency of each method. LANTERN++ consistently outperforms EAGLE-1 across all models, achieving higher step compression and lower latency. Notably, on Lumina-mGPT, LANTERN++ improves the step compression ratio from ×$\times 2.94$ to ×$\times 3.63$, reducing latency from ×$\times 2.10$ to ×$\times 2.56$. Similar trends are observed for Anole (×$\times 3.41$ vs. ×$\times 2.68$) and LlamaGen-XL Stage 2 (×$\times 2.86$ vs. ×$\times 2.38$). These results demonstrate that LANTERN++ enables deeper accepted sequences by decoupling token selection from low-confidence predictions, leading to greater acceleration.
In addition to efficiency, we assess the impact of LANTERN++ on image quality using Fréchet Inception Distance (FID) (Heusel et al., 2017) and CLIP score (Hessel et al., 2021). While LANTERN++ slightly increases FID due to its relaxed acceptance condition, the overall degradation remains minimal. For example, on Lumina-mGPT, FID increases from $28.93$ (standard AR) to $30.11$ ($\lambda=2$) and $33.91$ ($\lambda=3$), while maintaining a competitive CLIP score. This trade-off highlights the flexibility of LANTERN++ in balancing speed and quality by adjusting the multiplicative bound $\lambda$. To further assess visual fidelity, Figure 2 presents images generated by LANTERN++ vs. standard AR decoding on Lumina-mGPT. Despite the increased acceleration, the visual quality remains highly consistent, with no significant loss of detail or artifacts. This qualitative analysis aligns with the quantitative results, confirming that LANTERN++ maintains strong generation quality while substantially reducing inference time. More comprehensive experimental results are provided in Appendix C.

### v2

# LANTERN++: Enhancing Relaxed Speculative Decoding with Static Tree Drafting for Visual Auto-regressive Models

@@ Abstract
Speculative decoding has been widely used to accelerate auto-regressive (AR) text generation. However, its effectiveness for visual AR models remains limited due to token selection ambiguity, where multiple tokens share similarly low probabilities and thus reduce acceptance rates. Recently, relaxed speculative decoding with dynamic tree drafting was proposed to mitigate this ambiguity, demonstrating promising results in accelerating visual AR models. However, we observe that token selection ambiguity still negatively affects dynamic tree drafting, resulting in shallow draft trees and limited acceleration. To overcome this issue, we introduce LANTERN++, a refined framework that integrates static tree drafting with a tailored relaxed acceptance condition, allowing drafts to be selected independently of low-confidence predictions. This enables the acceptance of deeper sequences, improving decoding efficiency while preserving image quality. Extensive experiments on state-of-the-art visual AR models demonstrate that LANTERN++ significantly accelerates inference, achieving up to ×$\mathbf{\times 2.56}$ speedup over standard AR decoding while maintaining high image quality. The code is publicly available at https://github.com/jadohu/LANTERN.
@@ 1 Introduction
Recent advances in speculative decoding have significantly accelerated auto-regressive (AR) generation in language models (Leviathan et al., 2023; Cai et al., 2024; Li et al., 2024a; b). However, extending speculative decoding to visual AR models (Sun et al., 2024; Team, 2024; Chern et al., 2024; Liu et al., 2024) introduces unique challenges due to the inherent characteristics of image token distributions. In particular, Jang et al. (2025) identified token selection ambiguity as a major characteristic, where multiple next-tokens share similarly low probabilities. This ambiguity makes it difficult for the drafter model to accurately predict the next token, as even the target model itself lacks confident predictions. As a result, speculative decoding becomes less effective, with drafter predictions often inaccurate and thus frequently rejected by the target model.
Recently, LANTERN (Jang et al., 2025) was proposed to address this by introducing a relaxed acceptance condition leveraging latent space similarities. Specifically, it allows accepting drafter predictions when they are sufficiently close to the target predictions in latent space, even if they do not perfectly match, motivated by the observation that nearby tokens in latent space yield visually similar outcomes. While LANTERN successfully improved inference speed without compromising image quality, we find that token selection ambiguity still remains problematic within LANTERN’s dynamic tree drafting, thereby limiting its full potential. Our analysis reveals two key issues: (1) low drafter confidence results in shallow draft trees, preventing the formation of long accepted sequences, and (2) deterministic selection of draft tokens leads to overly low acceptance probabilities.
@@
To overcome these limitations, we introduce LANTERN++, an enhanced framework that revises LANTERN in two key aspects: (1) the adoption of static tree drafting and (2) the integration of a multiplicative bound tailored to static tree drafting. By employing static tree drafting, LANTERN++ ensures that deep draft sequences can be generated even under low-confidence predictions. Moreover, static tree drafting allows draft tokens to be selected stochastically rather than deterministically, effectively restoring a normal level of acceptance probability. Furthermore, LANTERN++ refines the relaxation mechanism by replacing the additive distributional distortion bound with a multiplicative bound, which scales acceptance probabilities relative to the original distribution. This adjustment prevents disproportionate favoring of certain tokens, ensuring stable probability scaling and preserving distributional consistency. Extensive experiments demonstrate that LANTERN++ achieves substantial speedups while maintaining high image quality, making it a promising direction for efficient visual AR modeling.
@@
Figure 1: Images generated by LANTERN++ on Lumina-mGPT (Liu et al., 2024) exhibit remarkable acceleration in decoding steps while maintaining high generation quality. The top row shows images generated by standard auto-regressive decoding, while the bottom row displays images generated with acceleration through LANTERN++. The step compression ratio is presented in white at the bottom right of each image.
@@
  • We identify critical limitations of dynamic tree drafting in visual AR models, showing that token selection ambiguity results in shallow draft trees and low acceptance rates, which significantly hamper speculative decoding performance.
  • We propose LANTERN++, a novel speculative decoding framework that refines LANTERN, which combines static tree drafting with a multiplicative relaxation mechanism, effectively unveiling the full potential of speculative decoding in visual AR models.
  • We demonstrate through extensive experiments that LANTERN++ achieves up to ×$\times 2.56$ reduction in latency and ×$\times 3.63$ step compression across multiple state-of-the-art visual AR models, while maintaining high image quality.
@@ 2 Preliminaries
2.1 Tree Drafting and Verification
@@
Unlike standard chain drafting (Leviathan et al., 2023), which generates a single draft token sequence, tree drafting (Cai et al., 2024; Li et al., 2024a; b) samples multiple draft tokens per forward pass, forming a tree structure for diverse draft sequences. Given an input $X_{1:N}$, the drafter first samples $s_{1}$ tokens $\hat{x}_{1,1},\ldots,\hat{x}_{1,s_{1}}$ from $p(x|X_{1:N})$, then expands each token with $s_{2}$ subsequent drafts in a single forward pass using a tree-aware attention mask. This process continues up to depth $K$, constructing a draft tree with only $K$ forward passes. The tree is then flattened and verified by the target model with an appropriate attention mask through a single forward pass.
@@
Tree drafting has two main variants: static tree drafting (Cai et al., 2024; Li et al., 2024a) and dynamic tree drafting (Li et al., 2024b). In static tree drafting, the draft tree structure is fixed, where a predefined number of draft tokens is always generated. In contrast, dynamic tree drafting adapts the tree structure based on the drafter model’s confidence. Specifically, it calculates a global accept probability $V_{i}$ for each token $t_{i}$, defined as:
$V_{i}=\prod_{t_{j}\in\text{Path}(\text{root},t_{i})}p_{j}\approx\prod_{t_{j}\in\text{Path}(\text{root},t_{i})}c_{j},$
where $p_{j}$ is the actual accept probability by the target model (unreachable during drafting) and $c_{j}$ is the drafter’s confidence score approximating $p_{j}$. When expanding the tree, the drafter model generates the top-$k$ candidate tokens for each current leaf node. It then trims the nodes based on their global accept probabilities $V_{i}$. As a result, dynamic tree drafting produces deeper or shallower draft trees depending on the drafter’s confidence at each step.
2.2 Token Selection Ambiguity and LANTERN
Visual AR models exhibit a phenomenon known as token selection ambiguity (Jang et al., 2025). In large language models (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023), the next-token prediction distribution typically forms a distinct, sharp peak, indicating high confidence in a single token choice. In contrast, the next-token distributions in visual AR models are considerably more dispersed. This is primarily due to the continuous, high-dimensional nature of visual tokens (e.g., pixels or image patches), which leads to more uncertain and diffuse next-token prediction distributions. Consequently, it becomes challenging for the drafter model to accurately match the target model’s top predictions, substantially limiting the potential of speculative decoding.
To mitigate this ambiguity, LANTERN (Jang et al., 2025) leverages the observation that tokens located closely in the latent space often represent visually similar and interchangeable semantics. Instead of requiring an exact match with the target model’s top prediction, LANTERN relaxes the acceptance condition by aggregating the target probabilities of tokens around the draft token. Intuitively, if a high-probability token predicted by the target model lies near the draft token, aggregating nearby token probabilities can increase the acceptance likelihood of the draft token. However, aggregating probabilities inherently distorts the original target distribution. To avoid excessive distributional shifts, LANTERN further introduces a constraint based on the Total Variation Distance (TVD).
Formally, for a given draft token $\widehat{x}$, LANTERN first identifies its $k$ nearest neighbors in latent space, denoted by $B_{k}(\widehat{x})$. Then, a TVD bound is applied, yielding a refined neighborhood set $A_{k,\delta}(\widehat{x})\subset B_{k}(\widehat{x})$ with $\delta>0$ such that the total amount of aggregated probability mass be less than $\delta$. The accept probability for $\widehat{x}$ is then relaxed as
[…]
If $\widehat{x}$ is not accepted, resampling is performed based on the adjusted probability distribution. This relaxation increases the effective acceptance rate while preserving visual fidelity, thereby alleviating the issues caused by token selection ambiguity. The more basic preliminaries on visual AR modeling and speculative decoding can be found in the Appendix A.
@@ 3 Why Dynamic Tree Drafting Fails in Visual AR Models
Applying dynamic tree drafting to visual AR models, where token selection ambiguity is prevalent, introduces two key issues: (1) draft trees become overly shallow due to low drafter confidence, and (2) the deterministic selection of draft tokens results in consistently low acceptance probabilities. These factors fundamentally limit the effectiveness of speculative decoding by restricting the acceptance of draft sequences, thereby reducing overall acceleration.
3.1 Shallow Draft Tree
As described in Section 2.1, dynamic tree drafting computes a global accept probability for each draft node in the draft tree as the product of the drafter’s confidence scores along the path from the root: $V_{i}=\prod_{t_{j}\in\text{Path}(\text{root},t_{i})}c_{j}$. However, in visual AR models, where drafter confidence scores tend to be low, this product decays exponentially as draft sequence length increases. Consequently, longer draft sequences become significantly less likely to survive, leaving only shorter ones. The result is a draft tree that is wide (many branches) but shallow, preventing the formation of long accepted sequences. Since longer sequences are key to achieving high step compression, this structural limitation restricts the acceleration potential of dynamic tree drafting. An example resulting shallow tree can be found in the Appendix C.1.
3.2 Low Accept Probabilities
In speculative sampling (Leviathan et al., 2023), a draft token $\widehat{x}$ is accepted with probability $\min(1,q(\widehat{x}|s)/p(\widehat{x}|s))$, which depends on the alignment between the drafter’s predictions and the target model’s likelihood. However, in dynamic tree drafting, tokens are deterministically selected based on the drafter’s confidence, effectively setting $p(\widehat{x}|s)=1$. This reduces the acceptance probability to $q(\widehat{x}|s)$. Due to token selection ambiguity, visual AR models produce highly dispersed next-token distributions, with even the most likely token often receiving a probability below 0.2. As a result, even accurate drafter predictions lead to an unnecessarily low acceptance rate.
@@ 3.2 Low Accept Probabilities
Table 1 highlights the contrasting effects of dynamic tree drafting in language models and visual AR models. In language models, dynamic tree drafting improves the step compression ratio from ×$\times 3.94$ to ×$\times 4.98$, leveraging concentrated next-token distributions and high-confidence drafter predictions to construct deeper draft trees. In visual AR models, however, the lower drafter confidence produces wide yet shallow draft trees and leads to low acceptance probabilities. Consequently, static tree drafting achieves a higher step compression ratio (×$\times 2.49$) than dynamic tree drafting (×$\times 2.11$), underscoring the latter’s limitations in handling token selection ambiguity.
@@ 4 LANTERN++: Relaxed Acceptance with Static Tree Drafting
As demonstrated in Section 3, dynamic tree drafting in visual AR models suffers from limited tree depth and low accept probabilities due to token selection ambiguity. Static tree drafting, in contrast, avoids these pitfalls by relying on a fixed tree structure rather than confidence-based adaptive expansion. However, it still struggles with low accept probabilities, as visual AR models inherently assign low probabilities to individual tokens. To address this, we introduce LANTERN++, a refined relaxation of the acceptance condition tailored for static tree drafting.
A key challenge in applying relaxation methods to static tree drafting is the variability of $p(x|s)$. Unlike dynamic tree drafting, where draft tokens are deterministically selected ($p(x|s)=1$), static tree drafting maintains the true drafter probabilities, which can vary significantly across tokens. This variability makes the original additive relaxation strategy problematic: when $p(x|s)$ is small, an additive increase $\delta$ can lead to disproportionately high accept probabilities, while for larger $p(x|s)$, the boost becomes negligible. This lack of proportionality leads to inconsistent behavior of the relaxation mechanism.
To overcome this issue, we replace the additive relaxation $\delta$ with a multiplicative distributional distortion bound $\lambda$, ensuring that accept probability adjustments remain proportional to the target model’s likelihood. Specifically, for each draft token $\widehat{x}$, its neighborhood $B_{k}(\widehat{x})$ is first determined in the same manner as in LANTERN. The selection of refined subset $A_{k,\lambda}(\widehat{x})\subseteq B_{k}(\widehat{x})$ is then changed such that
[…]
@@
This multiplicative relaxation offers two key advantages. First, it maintains a consistent proportional adjustment across different values of $p(\widehat{x}|s)$, ensuring that the acceptance probability is neither excessively inflated nor suppressed. Second, it prevents instability in cases when $p(\widehat{x}|s)$ is small, as a fixed additive boost could lead to over-amplification. By contrast, the multiplicative bound scales naturally with $q(x|s)$, yielding a more stable and controlled relaxation.
In summary, LANTERN++ improves upon LANTERN through two key components: (1) the use of static tree drafting, and (2) the introduction of a multiplicative relaxation $\lambda$. Static tree drafting enables the robust construction of deep draft trees even under severe token selection ambiguity, and the multiplicative bound further enhances this by providing a stable and consistent relaxation mechanism, complementing the static tree drafting.
@@ 5 Experiments
We conduct extensive experiments to evaluate the effectiveness of LANTERN++ in accelerating inference for visual AR models while preserving image quality. Our evaluation compares LANTERN++ against standard autoregressive (AR) decoding and EAGLE-1 (Li et al., 2024a) across multiple state-of-the-art visual AR models, including Lumina-mGPT (Liu et al., 2024), Anole (Chern et al., 2024), and LlamaGen-XL (Sun et al., 2024). All experiments are conducted on the MS-COCO 2017 validation set (Lin et al., 2014).
[…]
Table 2 presents the step compression ratios and inference latency of each method. LANTERN++ consistently outperforms EAGLE-1 across all models, achieving higher step compression and lower latency. Notably, on Lumina-mGPT, LANTERN++ improves the step compression ratio from ×$\times 2.94$ to ×$\times 3.63$, reducing latency from ×$\times 2.10$ to ×$\times 2.56$. Similar trends are observed for Anole (×$\times 3.41$ vs. ×$\times 2.68$) and LlamaGen-XL Stage 2 (×$\times 2.86$ vs. ×$\times 2.38$). These results demonstrate that LANTERN++ enables the acceptance of deeper sequences by decoupling token selection from low-confidence predictions, leading to greater acceleration.
In addition to efficiency, we evaluate the impact of LANTERN++ on image quality using the Fréchet Inception Distance (FID) (Heusel et al., 2017) and the CLIP score (Hessel et al., 2021). While LANTERN++ slightly increases FID due to its relaxed acceptance condition, the overall degradation remains minimal. For example, on Lumina-mGPT, FID increases from $28.93$ (standard AR) to $30.11$ ($\lambda=2$) and $33.91$ ($\lambda=3$), while maintaining a competitive CLIP score. This trade-off highlights the flexibility of LANTERN++ in balancing speed and quality by adjusting the multiplicative bound $\lambda$.
To further assess visual fidelity, Figure 2 presents images generated by LANTERN++ vs. standard AR decoding on Lumina-mGPT. Despite the increased acceleration, the visual quality remains highly consistent, with no significant loss of detail or artifacts. This qualitative analysis aligns with the quantitative results, confirming that LANTERN++ maintains strong generation quality while substantially reducing inference time. More comprehensive experimental results are provided in Appendix C.
Table 2: Acceleration performance and image generation performance of standard decoding, EAGLE-1, and LANTERN++ on various visual AR models using MS-COCO 2017 validation captions.